OpenStudy (anonymous):

write an equation in slope intercept form for the line that is parallel to the given line and that passes through the given point: 4y=x;(4,0)

OpenStudy (anonymous):

1. Find the slope of the current line: 4y=x first re-write the equation: y = x/4 compare with general line equation: y = mx + c, where m is the slope so current line: y = x/4 => slope is 1/4 A line perpendicular to this will have a slope of -4/1 therefore the equation = y = -4x + c now plug in the point above (4,0) => 0 =-4*4 +c => 16 =c Therefore the equation of the new line is: y=-4x +16

OpenStudy (anonymous):